OpenStudy (anonymous):

How do I simplify the difference quotient where (7)/(x+8)?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

You are suppose to plug it into the difference quotient formula

OpenStudy (anonymous):

try this :) http://www.youtube.com/watch?v=1O5NEI8UuHM

OpenStudy (anonymous):

what is basically is though:\[f(x)={7\over(x+8)}\]\[\text{difference quotient}:{f(x+h)-f(x)\over h}\]basically you just have to plug in x+h in place of x and then subtract f(x).. like this:\[\large {7\over (x+h+8)}-{7\over(x+8)}\over h\]

OpenStudy (anonymous):

From there where do you go? It confuses me I guess because it is a fraction.

OpenStudy (anonymous):

\[\large {7\over (x+h+8)}-{7\over(x+8)}\over h\]\[{7(x+8)-7(x+h+8)\over (x+8)(x+h+8)}\times {1\over h}\]simplify:\[{7x+56-7x-7h-56\over x^2+hx+8x+8x+8h+64}\times{1\over h}\]\[{-7h\over h(x^2+hx+16x+8h+64)}\]\[={-7\over x^2+hx+16x+8h+64}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem! :)